OpenStudy (anonymous):

MEDAL AND FAN Complete the standard form of the equation 3y/2 = 8 - x ?x = ?y = 16

OpenStudy (anonymous):

I get it! :p

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[\frac{ 3y }{ 2 } = 8 - x\] That is the question

OpenStudy (anonymous):

right

OpenStudy (anonymous):

\[\frac{ 3y }{ 2 } - 8 +x = 0\] Take LCM: \[\frac{ 2x + 3y - 16 }{ 2 } = 0\]

OpenStudy (anonymous):

move the 2 to the right hand side. \[2x-3y-16=0\] Move the (-16) to the right hand side as well: You get: \[2x-3y=16\]

OpenStudy (anonymous):

that's it :) Did you understand?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What is the x-intercept of the line with the equation 8x - 1/3y = 16 ( ?, 0 )

OpenStudy (anonymous):

In x intercept y=0, so put y=0 i.e 8x- 1/3(0)=16 8x-0 =16 8x=16 x=16/8 x=2

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

Which line is a graph of the equation: 2x + 5y = 10? http://static.k12.com/calms_media/media/213000_213500/213123/1/7c34fa0e2dcd286d023240417eb68498aae2b193/VHS_ALG_S1_07_L203_L303_LQ4_graphi.gif A. line a B. line b C. line c D. line d

OpenStudy (anonymous):

@Iffat0311

OpenStudy (anonymous):

srry

OpenStudy (anonymous):

I don't know this one, but it's either line b or d.

OpenStudy (anonymous):

Which point is collinear with points A and D? http://static.k12.com/calms_media/media/202500_203000/202592/1/8a992e88dbbabd6dcaf1dd91a8c102cfd382b597/28851_000515-23_ChoiceD.jpg A. (1, 1) B. (6, –8) C. (–4, 3) D. (1, –1)

OpenStudy (anonymous):

i can't see the question sorry :p and I gotta go. it's too late.

OpenStudy (anonymous):

k thx